A bearer bond worth Rs. 50,000 after six years. What will be the present value of the bond if interest is

added yearly at 12 percent.


1
Expert's answer
2021-04-20T07:36:32-0400

"PV = \frac{FV}{(1+r)^n}"


PV = present value

FV = future value

r = interest r/100

n = number of periods


"PV = \frac{ 50000}{[1+\frac{12}{100}]^6}=25,331.56"


Present value = Rs 25,331.56